# A causal inference framework for spatial confounding

### v7

@@ 8 Discussion
n method bias sd mse coverage
$1.000\times 10^{3}$ RSR $6.784\times 10^{-1}$ $3.030\times 10^{-2}$ $4.611\times 10^{-1}$ 0%
$1.000\times 10^{3}$ spline $2.763\times 10^{-1}$ $2.757\times 10^{-2}$ $\mathbf{7.712\times 10^{-2}}$ 0%
$1.000\times 10^{3}$ gp $3.755\times 10^{-1}$ $2.176\times 10^{-2}$ $1.415\times 10^{-1}$ 0%
$1.000\times 10^{3}$ gSEM $5.868\times 10^{-1}$ $3.014\times 10^{-1}$ $4.350\times 10^{-1}$ 73%
$1.000\times 10^{3}$ spatial+ $\mathbf{2.448\times 10^{-1}}$ $3.007\times 10^{-1}$ $1.502\times 10^{-1}$ 94%
$1.000\times 10^{3}$ svc_mle $3.640\times 10^{-1}$ $4.203\times 10^{-2}$ $1.342\times 10^{-1}$
$1.000\times 10^{3}$ spline_interaction $3.513\times 10^{-1}$ $6.681\times 10^{-2}$ $1.279\times 10^{-1}$ 99%
$1.000\times 10^{3}$ gp3 $5.565\times 10^{-1}$ $8.876\times 10^{-2}$ $3.175\times 10^{-1}$
$1.000\times 10^{3}$ spline3 $1.153\times 10^{0}$ $1.582\times 10^{-1}$ $1.353\times 10^{0}$
$1.000\times 10^{3}$ rf $5.394\times 10^{-1}$ $8.793\times 10^{-2}$ $2.986\times 10^{-1}$ 0%
$1.000\times 10^{3}$ BART $7.024\times 10^{-1}$ $1.265\times 10^{-1}$ $5.093\times 10^{-1}$ 1%
$1.000\times 10^{4}$ RSR $6.750\times 10^{-1}$ $9.106\times 10^{-3}$ $4.557\times 10^{-1}$ 0%
$1.000\times 10^{4}$ spline $2.267\times 10^{-1}$ $1.188\times 10^{-2}$ $5.152\times 10^{-2}$ 0%
$1.000\times 10^{4}$ gp $3.229\times 10^{-1}$ $7.812\times 10^{-3}$ $1.043\times 10^{-1}$ 0%
$1.000\times 10^{4}$ gSEM $2.386\times 10^{-1}$ $7.345\times 10^{-2}$ $6.231\times 10^{-2}$ 10%
$1.000\times 10^{4}$ spatial+ $\mathbf{2.049\times 10^{-1}}$ $7.285\times 10^{-2}$ $\mathbf{4.726\times 10^{-2}}$ 20%
$1.000\times 10^{4}$ spline_interaction $3.150\times 10^{-1}$ $4.915\times 10^{-2}$ $1.016\times 10^{-1}$ 98%
$1.000\times 10^{4}$ spline3 $1.292\times 10^{0}$ $9.115\times 10^{-2}$ $1.676\times 10^{0}$
$1.000\times 10^{4}$ rf $5.344\times 10^{-1}$ $3.208\times 10^{-2}$ $2.867\times 10^{-1}$ 0%
$1.000\times 10^{4}$ BART $5.844\times 10^{-1}$ $6.789\times 10^{-2}$ $3.461\times 10^{-1}$ 0%

### v8

@@ 8 Discussion
{adjustwidth}
-3cm-3cm n method bias sd mse coverage $1.000\times 10^{3}$ RSR $6.784\times 10^{-1}$ $3.030\times 10^{-2}$ $4.611\times 10^{-1}$ 0% $1.000\times 10^{3}$ spline $2.763\times 10^{-1}$ $2.757\times 10^{-2}$ $\mathbf{7.712\times 10^{-2}}$ 0% $1.000\times 10^{3}$ gp $3.755\times 10^{-1}$ $2.176\times 10^{-2}$ $1.415\times 10^{-1}$ 0% $1.000\times 10^{3}$ gSEM $5.868\times 10^{-1}$ $3.014\times 10^{-1}$ $4.350\times 10^{-1}$ 73% $1.000\times 10^{3}$ spatial+ $\mathbf{2.448\times 10^{-1}}$ $3.007\times 10^{-1}$ $1.502\times 10^{-1}$ 94% $1.000\times 10^{3}$ svc_mle $3.640\times 10^{-1}$ $4.203\times 10^{-2}$ $1.342\times 10^{-1}$ $1.000\times 10^{3}$ spline_interaction $3.513\times 10^{-1}$ $6.681\times 10^{-2}$ $1.279\times 10^{-1}$ 99% $1.000\times 10^{3}$ gp3 $5.565\times 10^{-1}$ $8.876\times 10^{-2}$ $3.175\times 10^{-1}$ $1.000\times 10^{3}$ spline3 $1.153\times 10^{0}$ $1.582\times 10^{-1}$ $1.353\times 10^{0}$ $1.000\times 10^{3}$ rf $5.394\times 10^{-1}$ $8.793\times 10^{-2}$ $2.986\times 10^{-1}$ 0% $1.000\times 10^{3}$ BART $7.024\times 10^{-1}$ $1.265\times 10^{-1}$ $5.093\times 10^{-1}$ 1% $1.000\times 10^{4}$ RSR $6.750\times 10^{-1}$ $9.106\times 10^{-3}$ $4.557\times 10^{-1}$ 0% $1.000\times 10^{4}$ spline $2.267\times 10^{-1}$ $1.188\times 10^{-2}$ $5.152\times 10^{-2}$ 0% $1.000\times 10^{4}$ gp $3.229\times 10^{-1}$ $7.812\times 10^{-3}$ $1.043\times 10^{-1}$ 0% $1.000\times 10^{4}$ gSEM $2.386\times 10^{-1}$ $7.345\times 10^{-2}$ $6.231\times 10^{-2}$ 10% $1.000\times 10^{4}$ spatial+ $\mathbf{2.049\times 10^{-1}}$ $7.285\times 10^{-2}$ $\mathbf{4.726\times 10^{-2}}$ 20% $1.000\times 10^{4}$ spline_interaction $3.150\times 10^{-1}$ $4.915\times 10^{-2}$ $1.016\times 10^{-1}$ 98% $1.000\times 10^{4}$ spline3 $1.292\times 10^{0}$ $9.115\times 10^{-2}$ $1.676\times 10^{0}$ $1.000\times 10^{4}$ rf $5.344\times 10^{-1}$ $3.208\times 10^{-2}$ $2.867\times 10^{-1}$ 0% $1.000\times 10^{4}$ BART $5.844\times 10^{-1}$ $6.789\times 10^{-2}$ $3.461\times 10^{-1}$ 0%